# Mingle: Mixture of Null-Space Gated Low-Rank Experts for Test-Time Continual Model Merging

### v2

@@ 1 Introduction
To address these limitations, recent works have explored an alternative paradigm known as continual model merging (CMM), which sequentially integrates independently fine-tuned models directly in parameter space, without revisiting any past training data [38, 51, 70]. CMM typically operates under a "merge-to-transfer" paradigm: given a pretrained model $\theta_{0}$ and independently fine-tuned models $\{\theta_{t}\}_{t=1}^{T}$, a unified model is constructed by combining task-specific weight updates $\Delta\theta_{t}=\theta_{t}-\theta_{0}$ via weighted averaging or projection-based strategies [24, 81, 69].
@@
In this paper, we propose Mingle (MIxtures of Null-Space Gated Low-rank Experts), a method designed to continually merge independently fine-tuned models at test-time while preserving prior knowledge. Mingle employs a mixture-of-experts architecture [27, 45] composed of lightweight LoRA-based [21] experts, enabling efficient and flexible test-time adaptation. To robustly prevent interference from previously learned tasks, we introduce a novel Null-Space Constrained Gating mechanism, restricting gating updates to task-orthogonal subspaces. Additionally, we propose an Adaptive Relaxation Strategy to dynamically modulate constraint strength based on test-time interference feedback during adaptation.
@@ 2 Related Work
Continual Learning. Continual learning (CL) seeks to mitigate catastrophic forgetting [43], where learning new tasks overwrites prior knowledge. Regularization-based methods constrain updates with importance weights [31, 92, 2, 30, 82], while distillation aligns outputs to preserve knowledge [20, 12, 60, 52]. Replay methods store exemplars or generate surrogates with prompts, prototypes, or generators [55, 39, 79, 61, 53], and dynamic architectures expand capacity via growth or ensembling [35, 98, 42]. Recent work leverages lightweight adapters or prompts in pre-trained models for efficient transfer [89, 23, 80]. Model merging offers an alternative route. Some methods remain close to conventional CL by sequentially fine-tuning and merging models to reduce forgetting [41, 42, 16], typically requiring training data. In contrast, continual model merging [28, 38, 5, 51, 70] merges independently fine-tuned models without revisiting prior data, enabling greater scalability and privacy.
@@ 3.1 Preliminaries
  • Continual Model Merging. A sequence of models $\{\theta_{i}\}$ are merged incrementally in parameter space without access to training data and earlier models: $\theta^{\text{merged}}_{t}=\text{Merge}(\theta^{\text{merged}}_{t-1},\theta_{t})$.
@@ 3.4 Adaptive Null-Space Constrained Gating for Interference Mitigation
Hard Null-Space Projection. After completing task $t$, we cache the $l$-th layer inputs in a seed buffer $\mathcal{D}_{t}^{\text{seed}}$ and estimate their covariance $\mathrm{Cov}_{t}^{(l)}\in\mathbb{R}^{d\times d}$. Applying truncated SVD yields the top-$k$ dominant subspaces $\tilde{U}_{t}^{(l)}\in\mathbb{R}^{d\times k}$. We then concatenate these with the subspaces from all previous tasks and orthonormalize: $U_{t}^{(l)}=\operatorname{orthonorm}\bigl[U_{t-1}^{(l)}|\tilde{U}_{t}^{(l)}\bigr]\in\mathbb{R}^{d\times tk}$. The hard projector is $P_{t}=I-U_{t}^{(l)}\,U_{t}^{(l)\top}\in\mathbb{R}^{d\times d}.$ To suppress interference from tasks $\leq t-1$, the gating update for task $t$ is $W_{t}^{(g,l)}\leftarrow W_{t}^{(g,l)}-\eta\,\nabla L_{t}^{(l)}P_{t-1}^{(l)},$. However, this projection may also discard gradient components that are informative for task $t$ whenever its feature support overlaps with $\operatorname{span}(U_{t-1})$.
@@ 4.2 Main Results
Overall Performance  As shown in Tab. 1, Mingle substantially outperforms prior continual merging methods on all CLIP backbones and task counts. It achieves the highest accuracy with backward forgetting kept near zero, demonstrating both strong forward learning and long-term stability. The lightweight variant performs on par with the full version, further underscoring the robustness of our approach. On NLP benchmarks (Tab. 2), Mingle likewise attains the best overall accuracy and non-negative BWT, improving on multiple GLUE tasks while maintaining balanced performance across the suite. Together, these results across vision and language confirm that Mingle consistently delivers state-of-the-art accuracy while nearly eliminating forgetting under diverse continual scenarios.
@@ 4.3 Ablation Results and Analysis
Ablation on TTA optimization steps Tab. 6 reports accuracy and backward transfer of CLIP ViT-B/32 under different adaptation steps across 8, 14, and 20 tasks. Accuracy improves steadily with longer schedules, rising from 60–63% at 5 steps to 77–86% at 50 steps. Forgetting remains negligible, with BWT close to zero in all cases and only a minor drop of about 2% in the 20-task setting at 50 steps. Notably, even 20 steps are sufficient to surpass all baselines in Tab. 1, while 50 steps provide the best accuracy at modest additional cost. These results confirm that Mingle is both effective under tight compute budgets and scalable with more adaptation.
@@
Number of Seed Samples The number of seed samples per class is crucial for TTA reliability and efficiency. As shown in Tab. 7 (c), using no samples reduces to the Static baseline, where LoRA modules are merged with fixed coefficients (0.3) without adaptation, yielding 70–79% accuracy. Introducing a single sample lifts accuracy to 81–88%, and adding more samples offers only minor gains. Variance across task orders decreases with more samples, making five samples per class a balanced trade-off between performance and efficiency.
@@
Hyper-parameter of Gate. We study the effect of three key hyper-parameters: $\gamma$, $\beta$, and $k$. $\gamma$ controls the strength of null-space suppression; smaller values lead to stronger attenuation of activations on prior tasks, reducing forgetting (Fig. 4(a)). As shown in Fig. 4(b), $\beta$ regulates the smoothness of the EMA used to accumulate interference signals. A moderate setting ($\beta=0.99$) balances responsiveness and stability. Smaller $\beta$ amplifies noise sensitivity, while larger $\beta$ slows detection of interference. $k$ determines the number of principal directions retained per task; performance saturates at $k=3$ (Fig. 4(c)), indicating that a small number of task-specific directions suffices.

### v3

@@ 1 Introduction
To address these limitations, recent works have explored an alternative paradigm known as continual model merging (CMM), which sequentially integrates independently fine-tuned models directly in parameter space, without revisiting any training data [38, 51, 70]. CMM typically operates under a "merge-to-transfer" paradigm: given a pretrained model $\theta_{0}$ and independently fine-tuned models $\{\theta_{t}\}_{t=1}^{T}$, a unified model is constructed sequentially by combining task-specific weight updates $\Delta\theta_{t}=\theta_{t}-\theta_{0}$ via weighted averaging or projection-based strategies [24, 81, 69].
@@
In this paper, we propose Mingle (MIxture of Null-Space Gated Low-Rank Experts), a method designed to continually merge independently fine-tuned models at test-time while preserving prior knowledge. Mingle employs a mixture-of-experts architecture [27, 45] composed of lightweight LoRA-based [21] experts, enabling efficient and flexible test-time adaptation. To robustly prevent interference from previously learned tasks, we introduce a novel Null-Space Constrained Gating mechanism, restricting gating updates to task-orthogonal subspaces. Additionally, we propose an Adaptive Relaxation Strategy to dynamically modulate constraint strength based on test-time interference feedback during adaptation.
@@ 2 Related Work
Continual Learning. Continual learning (CL) seeks to mitigate catastrophic forgetting [43], where learning new tasks overwrites prior knowledge. Regularization-based methods constrain updates with importance weights [31, 92, 2, 30, 82], while distillation aligns outputs to preserve knowledge [20, 12, 60, 52]. Replay methods store exemplars or generate surrogates with prompts, prototypes, or generators [55, 39, 79, 61, 53], and dynamic architectures expand capacity via growth or ensembling [35, 98, 42]. Recent work leverages lightweight adapters or prompts in pre-trained models for efficient transfer [89, 23, 80]. Model merging offers an alternative route. Some methods remain close to conventional CL by sequentially fine-tuning and merging models to reduce forgetting [41, 42, 16], typically requiring training data. In contrast, continual model merging [28, 38, 5, 51, 70] merges independently fine-tuned models without revisiting training data, enabling greater scalability and privacy.
@@ 3.1 Preliminaries
  • Continual Model Merging. A sequence of models are merged incrementally in parameter space without access to training data and earlier models: $\theta^{\text{merged}}_{t}=\text{Merge}(\theta^{\text{merged}}_{t-1},\theta_{t})$.
@@ 3.4 Adaptive Null-Space Constrained Gating for Interference Mitigation
Hard Null-Space Projection. After completing task $t$, we cache the $l$-th layer inputs in the seed buffer $\mathcal{D}_{t}^{\text{seed}}$ and estimate their covariance $\mathrm{Cov}_{t}^{(l)}\in\mathbb{R}^{d\times d}$. Applying truncated SVD yields the top-$k$ dominant subspaces $\tilde{U}_{t}^{(l)}\in\mathbb{R}^{d\times k}$. We then concatenate these with the subspaces from all previous tasks and orthonormalize: $U_{t}^{(l)}=\operatorname{orthonorm}\bigl[U_{t-1}^{(l)}|\tilde{U}_{t}^{(l)}\bigr]\in\mathbb{R}^{d\times tk}$. The hard projector is $P_{t}=I-U_{t}^{(l)}\,U_{t}^{(l)\top}\in\mathbb{R}^{d\times d}.$ To suppress interference from tasks $\leq t-1$, the gating update for task $t$ is $W_{t}^{(g,l)}\leftarrow W_{t}^{(g,l)}-\eta\,\nabla L_{t}^{(l)}P_{t-1}^{(l)}$. However, this projection may also discard gradient components that are informative for task $t$ whenever its feature support overlaps with $\operatorname{span}(U_{t-1})$.
@@ 4.2 Main Results
Overall Performance  As shown in Tab. 1, Mingle substantially outperforms previous continual merging methods on all CLIP backbones and task counts. It achieves the highest accuracy with backward forgetting kept near zero, demonstrating both strong forward learning and long-term stability. The lightweight variant performs on par with the full version, further underscoring the robustness of our approach. On NLP benchmarks (Tab. 2), Mingle likewise attains the best overall accuracy and non-negative BWT, improving on multiple GLUE tasks while maintaining balanced performance across the suite. Together, these results across vision and language confirm that Mingle consistently delivers state-of-the-art accuracy while nearly eliminating forgetting under diverse continual scenarios.
@@ 4.3 Ablation Results and Analysis
Ablation on TTA optimization steps. Tab. 6 reports accuracy and backward transfer of CLIP ViT-B/32 under different adaptation steps across 8, 14, and 20 tasks. Accuracy improves steadily with longer schedules, rising from 60–63% at 5 steps to 77–86% at 50 steps. Forgetting remains negligible, with BWT close to zero in all cases and only a minor drop of about 2% in the 20-task setting at 50 steps. Notably, as few as 20 steps are sufficient to surpass all baselines in Tab. 1, while 50 steps yield the best accuracy with only a modest increase in cost. These results confirm that Mingle is both effective under tight compute budgets and scalable with additional adaptation.
@@
Number of Seed Samples. The number of seed samples per class is crucial for TTA reliability and efficiency. As shown in Tab. 7 (c), using no samples reduces to the Static baseline, where LoRA modules are merged with fixed coefficients (0.3) without adaptation, yielding 70–79% accuracy. Introducing a single sample lifts accuracy to 81–88%, and adding more samples offers only minor gains. Variance across task orders decreases with more samples, making five samples per class a balanced trade-off between performance and efficiency.
@@
Hyper-parameter of Gate. We study the effect of three key hyper-parameters: $\gamma$, $\beta$, and $k$. $\gamma$ controls the strength of null-space suppression; smaller values lead to stronger attenuation of activations on prior tasks, reducing forgetting (Fig. 4(a)). As shown in Fig. 4(b), $\beta$ regulates the smoothness of the EMA used to accumulate interference signals. A moderate setting ($\beta=0.99$) balances responsiveness and stability. Smaller $\beta$ amplifies noise sensitivity, while larger $\beta$ slows detection of interference. $k$ determines the number of principal directions retained per task. The mitigation of forgetting saturates at $k=3$ (Fig. 4(c)), indicating that a small number of task-specific directions suffices.